given the sequence qn=(4n-3)/(2-5n), find q10 and q50.then find lim qn
Answer by venugopalramana(3286)

given the sequence qn=(4n-3)/(2-5n), find q10 and q50.then find lim qn
PUT N=10 TO GET Q10 AND N=50 TO GET Q50 IN THE ABOVE EQN.
Q10=(4*10-3)/(2-5*10)=-37/48=-0.771
Q50=(4*50-3)/(2-5*50)=-197/248=-0.794
AS REGARDS LIMIT QN , WE SHOULD KNOW LIMIT WHEN?
SUPPOSE LIMIT AS X GOES TO INFINITY IS
QN=(4-3/X)/(2/X-5)=-4/5=-0.8...IS THE LIMIT AS X TENDS TO INFINITY.
YOU CAN SE THIS FROM ABOVE ANSWERS.AT N=10,QN=-0.771...AT N=50,QN=-.794
LIKE THAT IT APPROACHES -0.8 AS N GROWS FURTHER LARGER AND LARGER
SUPPOSE LIMIT AS X GOES TO ZERO IS
QN=(4*0-3)/(2-5*0)=-3/2...IS THE LIMIT AS X TENDS TO ZERO.